The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                         http://www.astro.uiuc.edu/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 31 April 4, 1999

Happy holidays!

Contents

Abstracts of 10 refereed papers

Abstracts of Refereed Papers

Emission-Line Properties of the LMC Bubble N70

Brooke P. Skelton^{1}, William H. Waller^{2}, Richard F. Gelderman^{3}

Larry W. Brown^{4}, Bruce E. Woodgate^{4}, Adeline Caulet^{5}, & Robert A. Schommer^{6}

1 University of Washington, Astronomy Department, Seattle, Washington 98195, USA 2 Raytheon STX Corporation, Harvard-Smithsonian Center for Astrophysics, and Tufts University, Department of Physics and Astronomy, Robinson Hall, Medford MA 02138, USA 3 NASA Goddard Space Flight Center and Western Kentucky University, Department of Physics and Astronomy, Bowling Green, Kentucky 42101-3576, USA 4 NASA Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA 5 University of Illinois, Astronomy Department, Urbana, IL 61801, USA 6 National Optical Astronomy Observatories, Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile

We present a spectrophotometric imaging study of the emission bubble N70 (DEM 301) in the Large Magellanic Cloud. N70 is approximately 100 pc in size with a nearly circular shell-like morphology.

The nebular emission is powered by an uncertain combination of EUV photons, intense winds, and supernova shock waves from the central population of high-mass stars (the OB association LH 114). We have obtained narrow-band images (FWHM ∼ 6 Å) of N70 in the light of Hαλ6563, [N II]λ6584, [S II]λλ6717,6731, and [O III]λ5007, along with the corresponding red and green continua. The resulting line fluxes and flux ratios are used to derive ionization rates, nebular densities, volume filling fractions, and excitation indices. The photoionizing luminosity inferred from the embedded stellar population is more than adequate to account for the observed hydrogen ionization rate.

We compare the emission-line photometry with that derived from similar imaging of the Orion nebula and with data collected from the literature on other emission-line regions in the LMC. Compared to the Orion nebula, N70 shows much higher [S II]/Hα intensity ratios which increase smoothly with radius — from <0.3 near the center to >1.0 towards the outer filamentary shell. The measured intensity ratios in N70 more closely match the range of excitation spanned by giant and supergiant H II shells and by some of the supernova remnants observed in the LMC.

The contending ionization and excitation processes in the interior and outer shell of N70 are evaluated in terms of the available data. EUV photons probably contribute most of the inner nebula’s ionization, whereas a combination of photoionization plus collisional ionization and excitation of sulfur atoms by low-velocity shocks seems to best fit the emission-line luminosities and intensity ratios observed in the outer shell. Considerations of the radiative and mechanical energetics that are involved may indicate the need for one or two supernova explosions having occurred during the last ∼Myr.

Accepted by: Publications of the Astronomical Society of the Pacific

For preprints, contact waller@cosmos2.phy.tufts.edu

Also available from the URL astro-ph/9903001

The structure and dynamics of the unique Honeycomb Nebula in the halo of 30 Doradus

M.P. Redman^{1}, Z.A. Al-Mostafa^{1,2}, J. Meaburn^{1}, M. Bryce^{1}, and J.E. Dyson^{3}

1 Department of Physics & Astronomy, University of Manchester, Oxford Road, Manchester M13 9PL, UK 2 King Abdulaziz City for Science and Technology, Institute of Astronomical and Geophysical Research, PO Box 6086, Riyadh 11442, Saudi Arabia 3 Department of Physics & Astronomy, University of Leeds, Leeds, LS2 9JT, UK

The structure and dynamics of the Honeycomb nebula in the Large Magellanic cloud are modelled as resulting from the interaction of a supernova explosion with an existing giant shell of the type commonly found in the LMC. In particular, the processes that may have led to the striking morphology and velocity structure of this nebula are examined. New data of the Honeycomb nebula, are presented. Here, spatially resolved profiles of the [S ii] 6716 & 6731Å density sensitive lines, obtained with the Manchester Echelle Spectrometer, supplement previous, complementary, Hα and [N ii] 6584Å profiles.

The agreements and differences between all of these data and this simple model are highlighted.

Accepted by: Astronomy and Astrophysics For preprints, contact mpr@ast.man.ac.uk

Also available from the URL ftp://www.ast.man.ac.uk/outgoing/mpr/honey.ps.gz

HST Observations of the very young SMC “blob” N 88A

M. Heydari-Malayeri^{1}, V. Charmandaris^{1}, L. Deharveng^{2}, M.R. Rosa^{3}, and H. Zinnecker^{4}

1 demirm, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 2 Observatoire de Marseille, 2 Place Le Verrier, F-13248 Marseille Cedex 4, France 3 Space Telescope European Coordinating Facility, European Southern Observatory, Karl-Schwarzschild-Strasse-2, D-85748 Garching bei München, Germany 4 Astrophysikalisches Institut Potsdam, An der Sternwarte 16, D-14482 Potsdam, Germany

High-resolution Hubble Space Telescope images have allowed us for the first time to resolve the compact SMC ionized “blob” N 88A (diameter ∼ 3′′.5 or 1 pc). This very young H ii region, which is hatching from its natal molecular cloud, is heavily affected by absorbing dust associated with the cloud. The interstellar reddening towards N 88A is on average A_{V} ∼ 1.5 mag and strikingly rises to more than 3.5 mag in a narrow dust band crossing the core of the H ii region. Such a high extinction is unprecedented for an H ii region in the metal-poor SMC. We present the photometry of some 70 stars lying towards the OB association at the center of which lies N 88A. The exciting star(s) of N 88A is not detected, due to the heavy extinction. The chronology of star formation is discussed for the whole region.

Submitted to: Astronony & Astrophysics For preprints, contact heydari@obspm.fr

UV Spectral Classifications of O and B Stars in the Large Magellanic Cloud

Margaret M. Smith Neubig^{1} and Frederick C. Bruhweiler^{1}

1 Institute for Astrophysics and Computational Sciences, Department of Physics, 200 Hannan Hall, The Catholic University of America, Washington, DC 20064, USA

In this extension of our previous ultraviolet classification study of Small Magellanic Cloud (SMC) stars, we construct an ultraviolet classification system for stars of the Large Magellanic Cloud (LMC), a galaxy with a metallicity intermediate to those of the Milky Way and the SMC. Employing the same method as was used for the SMC, we compile a set of LMC spectra drawn from the International Ultraviolet Explorer (IUE) archives, and use these spectra to construct a framework of reliable UV spectral classification criteria. We classify 143 LMC IUE spectra, and identify the LMC standard stars which define our UV classification system. Representative spectra are presented to illustrate the patterns that the UV spectral classification features exhibit. The LMC and SMC UV classification systems are intended to be applicable to UV spectra of stars in external galaxies for which no MK standards have been defined. Such UV spectra are expected to be collected from spaceborne instruments such as the Space Telescope Imaging Spectrograph (STIS), aboard the Hubble Space Telescope.

Accepted by: The Astronomical Journal

For preprints, contact neubig@iue.gsfc.nasa.gov

The Fourth Catalogue of Population I Wolf-Rayet Stars in the Large Magellanic Cloud

J. Breysacher^{1}, M. Azzopardi^{2}, G. Testor^{3}

1 European Southern Observatory, Karl-Schwarzschild-Strasse 2, D-85748 Garching bei München, Germany 2 Observatoire de Marseille, 2 Place Le Verrier, F-13248 Marseille Cedex 04, France 3 Observatoire de Paris, section de Meudon, DAEC, F-92195 Meudon Cedex, France

The catalogue provides for each of the 134 W-R stars of Population I presently known in the Large Magellanic Cloud, accurate equatorial coordinates, photometric data, spectral classification, binary status, correlation with OB associations and HII regions. The miscellaneous designations of the stars are also listed. Although completeness is not pretended, results published during the last decade are highlighted in the notes given for each individual star. A uniform set of finding charts is presented.

Accepted by: Astronomy and Astrophysics Supplements For preprints, contact jbreysac@eso.org

Differences in the fractions of Be stars in galaxies

André Maeder^{1}, Eva K. Grebel^{2,3}, and Jean–Claude Mermilliod^{4}

1 Observatoire de Genève, CH-1290 Sauverny, Switzerland 2 UCO/Lick Observatory, University of California, Santa Cruz, CA 95064, USA 3 University of Washington, Department of Astronomy, Box 351580, Seattle, WA 98195, USA 4 Institut d’Astronomie, UniL, CH–1290 Chavannes–des–Bois, Switzerland

The number ratios Be/(B + Be) of Be to B–type stars in young, well studied clusters of the Galaxy, the LMC and SMC are examined. In order to disentangle age and metallicity effects we choose clusters in the same age interval and for which reliable photometric and spectroscopic data are available. Number counts are made for various magnitude intervals, and the results are found to be stable with respect to this choice. In the magnitude interval M_{V} = −5 to -1.4 (i.e. O9 to B3) we obtained a ratio Be/(B + Be) = 0.11, 0.19, 0.23, 0.39 for 21 clusters located in the interior of the Galaxy, the exterior of the Galaxy, the LMC and the SMC, respectively.

Various hypotheses for these differences are examined. An interesting possibility is that the average rotation is faster at low metallicities as a result of star formation processes. The much higher relative N–enrichment found by Venn (1999) in A–type supergiants of the SMC, compared to galactic supergiants, also strongly supports the presence of more rotational mixing at low metallicities. We discuss whether high rotational mixing may be the source of primary nitrogen in the early chemical evolution of galaxies.

Accepted by: Astronomy & Astrophysics

For preprints, contact Andre.Maeder@obs.unige.ch
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9904008

V(RI)_{C} Photometry of Cepheids in the Magellanic Clouds

Thomas G. Barnes III^{1}, Inese I. Ivans, Jodie R. Martin^{2}, C.S. Froning ^{2}, and Thomas J. Moffett^{3}

1 McDonald Observatory, The University of Texas at Austin, Austin, TX 78712-1083, USA 2 Department of Astronomy, The University of Texas, Austin, Texas 78712-1083, USA 3 Purdue University, Department of Physics, 1396 Physics Building, West Lafayette, Indiana 47907- 1396, USA

We present V(RI)_{C} data for thirteen Cepheids in the Large Magellanic Cloud and eight in the Small Magellanic Cloud. The total number of new measures is fifty-five in each wavelength band. The median uncertainty in the photometry is ±0.03 mag. These results supplement a larger photometric program on Magellanic Cloud Cepheids previously published.

Submitted to: The Astrophysical Journal Supplement Series For preprints, contact Thomas G. Barnes III tgb@astro.as.utexas.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9903095

Statistics of Stellar Populations of Star Clusters and Surrounding Fields in the Outer Disk of the Large Magellanic Cloud

J.F.C. Santos Jr.^{1}, A.E. Piatti^{2}, J.J. Clariá^{2}, E. Bica^{3}, D. Geisler^{4}, H. Dottori^{3}

1 Dep. de Física, ICEx-UFMG, C.P. 702, 30123-970 Belo Horizonte MG, Brazil 2 Observatorio Astronómico de Córdoba, Laprida 854, 5000, Córdoba, Argentina 3 Dep. de Astronomia, IF-UFRGS, C.P. 15051, 91501-970 Porto Alegre RS, Brazil 4 Dep. de Física, FCFM, Universidad de Concepción, Casilla 160-C, Concepción, Chile

A comparative analysis of Washington color-magnitude diagrams (CMDs) for 14 star clusters and respective surrounding fields in the Large Magellanic Cloud (LMC) outer disk is presented. Each CCD frame including field and respective cluster covers an area of 185□′. The stellar population sampled is of intermediate age and metallicity. CMD radial analysis involving star count ratios, morphology and integrated light properties are carried out. Luminosity functions (LFs) are also presented. Two main results are: (i) Within the range 4 < R(kpc) < 8, the distance from the LMC center is well correlated with the average age in the sense that inner fields are younger and; (ii) Beyond ≈8kpc the outer fields do not show evidence of a significant intermediate-age component in their stellar populations, as inferred from red giant clump star counts.

Accepted by: The Astronomical Journal

For preprints, contact jsantos@fisica.ufmg.br
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9903329

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. I. Double-Mode Cepheids in the Small Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1} P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present a sample of 93 double-mode Cepheids detected in the 2.4 square degree area in the central part of the SMC. 23 stars from the sample pulsate in the fundamental mode and the first overtone while 70 objects are the first and second overtone pulsators. This is the largest sample of such type Cepheids detected in one environment so far.

We analyze period ratio of double-mode Cepheids and Fourier parameters of decomposition of the light curves of these objects. We also present location of different type Cepheids from the SMC in the color-magnitude diagram and show their distribution of V − I color indices. We find one object which is probably a blend, either physical or optical, of two Cepheids pulsating in the fundamental mode.

Accepted by: Acta Astronomica 49, 1.

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9903391

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. II. Single-Mode Second Overtone Cepheids in the Small Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1} P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present a sample of 13 candidates for single-mode second overtone Cepheids discovered in the 2.4 square degree area of the SMC bar during the OGLE microlensing search. All candidates passed four photometric tests: analysis of Fourier parameters of the light curve, R_{21} and ϕ_{21}, position in the period-luminosity diagram and location in the color-magnitude diagram.

The stars form a very homogeneous group of objects with very similar light curves of low amplitude and almost sinusoidal shape. Their color indices B − V and V − I are also very similar and the objects determine the observational blue edge of the instability strip of the first overtone Cepheids. This is the first firm sample of the single-mode second overtone Cepheids detected so far.

Accepted by: Acta Astronomica 49, 45.

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9903393